OpenStudy (anonymous):

If a bacteria population starts with 150 bacteria and doubles every four hours, then the number of bacteria after t hours is n = f(t) = 150 · 2t/4. (a) Find the inverse of this function.

OpenStudy (anonymous):

please help!!!1

OpenStudy (anonymous):

Just express t in terms of f(x), thus: t = 4f(t)/150 You could then calculate after how many hours the population was, say, 1050. Answer t = 1050 times 4/150 = 28 hours

OpenStudy (anonymous):

can you show me how you got this please

OpenStudy (anonymous):

i think it is \[150\times 2^{\frac{t}{4}}\]

OpenStudy (anonymous):

so inverse is \[\frac{4ln(\frac{x}{150})}{ln(2)}\]